OpenStudy (3psilon):

How to find sec(7(pi)/6) I forgot how to do this with a calculator and without a calculator . Help please :)

OpenStudy (anonymous):

\[\sec x = \frac{ 1 }{ \cos x }\]

OpenStudy (3psilon):

so \[1/ -\sqrt{3}/2?\] @Shi

OpenStudy (anonymous):

yup you're correct, sorry openstudy stopped loading for me